OpenStudy (anonymous):

Write the equation y - 2 = 2(x - 3) in standard form.

OpenStudy (anonymous):

y-2=2x-6 y=2x-6+2 y=2x-4

OpenStudy (anonymous):

y - 2 = 2(x-3) y - 2 = 2x - 6 [expand the brackets] y= 2x - 4 [add +2 on each side] Slope = 2 Y intercept= -4 [Using y=mx+c, m being slope, c being y intercept]

OpenStudy (anonymous):

First, change it into Slope-Intercept form ~ y - 2 = 2(x - 3) Distribute. y - 2 = 2x - 6 Add 2. y = 2x - 4 Change into Standard form ~ y = 2x - 4 Subtract 2x from both sides. -2x + y = -4 No negatives aloud as A (ax + by = c). Divide by -1. 2x - y = 4

OpenStudy (anonymous):

So 2x - y = 4 is the standard form.

OpenStudy (anonymous):

y - 2x = 4 2x - y = -6 2x + y = 6 2x - y = 4

OpenStudy (anonymous):

oops sorry

OpenStudy (anonymous):

(: